How do we solve problems using the principles together such as -9/2+4=91/2
Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

How do we solve problems using the principles together such as -9/2+4=91/2
----------------
(-9/2) + 4
---
(-9/2) + (8/2)
---
= -1/2
========================
Cheers,
Stan H.

Answer by Theo(13342)

-9/2 + 4 equals -9/2 + 8/2 which equals (-9+8)/2 which equals -1/2.

+9/2 + 4 equals 9/2 + 8/2 which equals (9+8)/2 which equals 17/2 which equals 8 and 1/2.

9 and 1/2 would be equal to 18/2 + 1/2 which would be equal to (18+1)/2 which equals 19/2.

i'm not sure what principles you are talking about.

there are several arithmetic principles that could be involved.

are you dealing with proper fractions and improper fractions?

are you dealing with addition and subtraction of fraction?

give me a better clue as to what you are looking for and i should be able to help you.